OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5) A. 23 B. –53 C. –9 D. 3

OpenStudy (mertsj):

Multiply the two functions and replace x with -5 in the result

OpenStudy (anonymous):

how u do that

OpenStudy (mertsj):

The functions are binomials. Do you know how to multiply binomials?

OpenStudy (anonymous):

subtract four from both side

OpenStudy (mertsj):

(a+b)(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

ok i have no idea what ur talking bout thxs for the help though

OpenStudy (mertsj):

f(x) times g(x)= (-2x-7) times (-4x+3)

OpenStudy (mertsj):

And I have no idea what you are talking about when your problem asks you to multiply 2 things and you start talking about subtracting 4 from both sides.

OpenStudy (mertsj):

f(x)*g(x)=(-2x-7)(-4x+3)

OpenStudy (anonymous):

would that turn into 8x-6x+28x-21

OpenStudy (mertsj):

-2x(-4x)=8x^2

OpenStudy (anonymous):

so then it will be 8x^2-6x+28x-21. so would i sub all the x's with -5

OpenStudy (mertsj):

Combine -6x +28x = 22x

OpenStudy (mertsj):

Now replace all the x's with -5

OpenStudy (anonymous):

so it would be 8x^2+22x-21 so then it will be 200-110-21. so wouldnt u square the x -5*-5 =25*8=200 then 28*-5=110 then -21

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

\[8(-5)^2+22(-5)-21=200-110-21=90-21=69\]

OpenStudy (anonymous):

ok but 69 is not one of my chosies

OpenStudy (mertsj):

Then you might have copied something wrong. Double check.

OpenStudy (mertsj):

That is f times g, isn't it?

OpenStudy (mertsj):

Or is it f composite g?

OpenStudy (mertsj):

If it is f composite g it would be easiest to find g(-5) and substitute that answer into f.

OpenStudy (mertsj):

g(-5)=-4(-5)+3=23 f(23)=-2(23)-7=-46-7=-53

OpenStudy (anonymous):

well ima try that and see if i get it right thank you for all your help i wrote it just the way i got it